For which of these reactions will the difference betweenΔH° and ΔE° be the greatest?

a. 2H2O2(l) →2H2O(l) + O2(g) b. CaCO3(s)→ CaO(s) +CO2(g) c. NO(g) +O3(g) → NO2(g) + O2(g) d. 2C2H6(g) +7O2(g) → 4CO2(g) + 6H2O(l) e. 4NH3(g)+ 5O2(g) → 4NO(g) + 6H2O(g)
Answer #1
From elementary thermodynamic relation
H = E + pV you find the difference between ΔH° andΔE°
ΔH° - ΔE° = Δ(p·V)
Since the volume occupied by liquid and solid is small compared togas, it is sufficient to calculate Δ(p·V) for the gaseouscomponents. Treat the gases as ideal gases:
p·V = N·R·T
hence at standard conditions
Δ(p·V) = R·T°·ΔN

The change of total number of moles of gaseous components due toreaction is :
a) +1
b) +1
c) 0
d) -5
e) +1
So reaction has the greatest magnitude ofΔ(p·V) and therefore the greatest difference betweenΔH° and ΔE°.